A company has two investment possibilities, with the following cash inflows:

Investment Year 1 Year 2 Year 3
A $1,200 1,400 1,700
B $1,600 1,600 1,600

If the firm can earn 6 percent in other investments, what is the present value of investments A and B? Use Appendix B and Appendix D to answer the question. Round your answers to the nearest dollar.
PV(Investment A): $  
PV(Investment B): $  

If each investment costs $4,000, is the present value of each investment greater than the cost of the investment?
The present value of investment A is less than or greater than Item 3 the cost?
The present value of investment B is less than or greater than Item 4 the cost?
Answer #1

1)

Investment A:

PV(Investment A) = 1200 / (1 + 0.06)1 + 1400 / (1 + 0.06)2 + 1700 / (1 + 0.06)3

PV(Investment A) = 1,132.0755 + 1,245.995502 + 1,42.35278

PV(Investment A) = $3,805

Investment B:

PV(Investment B) = 1,600 / (1 + 0.06)1 + 1,600 / (1 + 0.06)2 + 1,600 / (1 + 0.06)3

PV(Investment B) = 1,509.434 + 1,423.9943 + 1,343.9085

PV(Investment B) = $4,277

2)

present value of investment A is less than investment

present value of investment B is greater than investment